OpenStudy (anonymous):

how to Find a cubic function with the given zeros square root 6, square root -6, -3

OpenStudy (kagıtucak):

(x-root 6)*(x+root 6)*(x+3) When you do the calculation you can get a cubic function. It is just a simple idea nad always works.

OpenStudy (anonymous):

how do you do the calculations

OpenStudy (aum):

If 'a' is a zero, then (x-a) is a factor. Given three zeros, the factors are: \((x+\sqrt{6})(x-\sqrt{6})(x+3) \) Multiply the first two factors. Make use of the identity: \((x+a)(x-a) = x^2 - a^2 \) Then multiply the result by (x+3).

OpenStudy (anonymous):

i understand thank you!!

OpenStudy (aum):

You are welcome.